# Time Weaver: A Conditional Time Series Generation Model

### v1

@@ D.2 Time Weaver-SSSD
Here, $A,B,C,D$ are transition matrices that are learned. Gu et al. (2022) propose stacking several SSM blocks together to create a Structured State Space sequence model (S4). Then, these SSM blocks are connected with normalization layers and point-wise FC layers in a way that resembles the transformer architecture. This architectural change is done to capture long-term dependencies in time series data. Alcaraz & Strodthoff (2023) adjusts this architecture to take label input, a binary vector of length 71. As shown in Fig. 8, we replaced this label input with the metadata embeddings obtained with our metadata preprocessing block to incorporate more various metadata modalities. We saw that this generates the best quality examples, and the remaining architecture is kept the same.
For the diffusion process, our experiments with Time Weaver-SSSD used 200 diffusion steps with the noise variance schedule values of $\beta_{1}=0.0001$ and $\beta_{T}=0.02$
Now, we provide the list of design choices and hyperparameter choices used in the Time Weaver-SSSD model.
@@
Table 7: Hyperparameters for Time Weaver-SSSD Architecture.
Appendix E GAN baselines
E.1 Main GAN baselines
For our main GAN baselines, we used Pulse2PulseGAN (Thambawita et al., 2021) and WaveGAN (Donahue et al., 2019). Since these approaches are not fundamentally conditional, we added additional layers to enable conditional generation.
  • For the Electricity and the ECG datasets, we used the implementation provided by (Thambawita et al., 2021) and (Alcaraz & Strodthoff, 2023). Since these datasets only have categorical metadata, we represented each categorical label by a fixed embedding. This fixed embedding was added to the output of each layer in the generator after the batch normalization layers. Similarly, we added the fixed embedding to the output of each layer in the discriminator. To learn the conditional distribution, along with predicting whether a sample is real or fake, we also predicted the logit of each categorical metadata, similar to (Odena et al., 2017). In our experiments, we noticed that predicting the metadata category for the fake sample rarely helps and provides poor-quality samples. Hence, we only predicted the category for the real samples.
@@ E.1 Main GAN baselines
E.1.1 WaveGAN implementation details
We trained the WaveGAN model for all the datasets for 1500 epochs with a learning rate of $10^{-4}$ and stored the checkpoints after every 100 epochs. We sampled noise, a $48$ dimensional vector for the Electricity, Air Quality, and Traffic datasets, and a $100$ dimensional vector for ECG. We relied on the pytorch implementation (Link to the repo) of WaveGAN and (Alcaraz & Strodthoff, 2023) for our experiments. We adjusted the number of parameters in the generator and discriminator to roughly match the Time Weaver models.
  • For the Air Quality dataset, the total number of trainable parameters in the GAN model is 15.2 million and the generator has 8.51 million trainable parameters.
  • For the Traffic dataset, the total number of trainable parameters in the GAN model is 13.7 million and the generator has 7.017 million trainable parameters.
  • For the Electricity dataset, the total number of trainable parameters in the GAN model is 13.3 million and the generator had 7.17 million parameters.
@@ E.1.1 WaveGAN implementation details
E.1.2 Pulse2PulseGAN implementation details
We trained the Pulse2PulseGAN model in the same manner as the WaveGAN for all the datasets. We trained the Pulse2PulseGAN model for 1500 epochs with a learning rate of $10^{-4}$ and stored the checkpoints after every 100 epochs. Here, the noise input to the generator had the same dimensions as the time series sample that we wanted to generate. We adjusted the number of parameters in the generator and discriminator to roughly match the Time Weavermodels.
  • For the Air Quality dataset and the Traffic, the total number of trainable parameters in the GAN model is 14.1 million and the generator has 7.45 million trainable parameters.
  • For the Electricity dataset, the total number of trainable parameters in the GAN model is 16.9 million and the generator has 8.4 million parameters.
@@ E.1.2 Pulse2PulseGAN implementation details
E.2 Additional GAN baselines
In addition to WaveGAN (Donahue et al., 2019) and Pulse2Pulse (Thambawita et al., 2021) models, we have implemented TTS-GAN (Li et al., 2022) and well-established TimeGAN (Yoon et al., 2019) method. Unfortunately, we were unable to train these models to generate effectively. These models were likely challenged by higher input lengths than their original implementation, where TimeGAN and TTS-GAN consider time steps up to 24 and 188, respectively, while we consider time steps of up to 1000. A similar problem was also faced in literature (Alcaraz & Strodthoff, 2023). We include our training examples after 10000 epochs for the traffic dataset in Figure 9.
Figure 9: TimeGAN and TTS-GAN failed to generate realistic samples after 10000 epochs This figure shows the samples generated for the test examples after 100, 1000, and 10000 training epochs, where row 1 and row 2 correspond to TTS-GAN and TimeGAN respectively. We can clearly see that both models fail to generate high-quality realistic samples.
Appendix F Evaluation Metrics
@@ F.1 J-FTSD details
For the Electricity, Air Quality, and Traffic datasets, the horizon is 96, i.e., $L=96$. So, we took time series and metadata patches of length $L_{\textnormal{patch}}=64$ from time step 1, i.e., 1 to 64, 2 to 65, …, and obtained the time series and the metadata embeddings using $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ respectively. We computed the J-FTSD from these embeddings using Eq. 6. For the ECG dataset, since the horizon is 1000, and the patch length is 256, we sampled patches of length 256 after every 10 time steps, i.e., 1 to 256, 10 to 266, etc.
One of the key points to be noted is that the feature extractors, $\phi_{\textnormal{time}}$, and $\phi_{\textnormal{meta}}$, are trained on the entire data distribution. This was done to ensure that the inefficiency of the feature extractors to extract accurate and metadata-specific features should not affect the evaluation process, which can occur when the feature extractors are trained on a training split alone. Therefore, we trained the feature extractors on the entire dataset to evaluate generative models.
F.2 Train on Synthetic Test on Real details (TSTR)
@@
For TSTR, we use a standard ResNet 1D (He et al., 2016) architecture. We performed the following classification tasks:
@@
  • Classification over months in the Electricity dataset. There are 12 classes in total and we trained the classifier with cross-entropy loss for 500 epochs with a learning rate of $10^{-4}$.
@@
  • Classification over heart disease statements in the ECG dataset. There are 71 classes and for a given time series sample, more than one class could be active. So, we trained a classifier with binary cross-entropy loss for 500 epochs with a learning rate of $10^{-4}$.
@@
  • Classification over the coarse weather description in the Traffic dataset. 11 coarse weather descriptions are available as annotations for each time step in the traffic dataset. To this end, we treated the classification task here as a multi-class, multi-label classification problem. So, we trained a classifier with binary cross-entropy loss for 200 epochs with a learning rate of $10^{-4}$.
@@
  • Classification over 12 weather stations for the Air Quality dataset. We used the cross-entropy loss for 200 epochs with a learning rate of $10^{-4}$.
@@
Here, we note that with the trained diffusion model, we generated the synthetic train, validation, and test datasets. The classifier is trained on the synthetic train dataset, and the checkpoints are stored with the synthetic validation dataset. We finally evaluated the model on the real test dataset.
Appendix G Additional Qualitative Results
@@
Figure 10: Time Weaver-CSDI Qualitative Results for Traffic Dataset
Figure 11: Time Weaver-CSDI Qualitative Results for the Air Quality Dataset
@@
Figure 14: Generated time series samples from the Electricity Dataset

### v2

@@ A.4.2 Time Weaver-SSSD
Here, $A,B,C,$ and $D$ are learnable transition matrices. Gu et al. (2022) propose stacking several SSM blocks together to create a Structured State Space sequence model (S4). These SSM blocks are connected with normalization layers and point-wise FC layers in a way that resembles the transformer architecture. This architectural change is done to capture long-term dependencies in time series data. Alcaraz & Strodthoff (2023) adjust this architecture to take label input, a binary vector of length 71. As shown in Fig. 8, we replaced this label input with the metadata embeddings obtained from our metadata preprocessing block to incorporate more complex metadata modalities.
For the diffusion process, our experiments with Time Weaver-SSSD used 200 diffusion steps with the noise variance schedule values of $\beta_{1}=0.0001$ and $\beta_{T}=0.02$.
Now, we provide the list of design choices and hyperparameters used in the Time Weaver-SSSD model in Table 7.
@@
Table 7: Hyperparameters for the Time Weaver-SSSD architecture.
A.5 GAN Baselines
A.5.1 Main GAN Baselines
For our main GAN baselines, we use Pulse2Pulse GAN (Thambawita et al., 2021) and WaveGAN (Donahue et al., 2019). Since these approaches are not fundamentally conditional, we added additional layers to enable conditional generation.
  • For the Electricity and ECG datasets, we used the implementation provided by (Thambawita et al., 2021) and (Alcaraz & Strodthoff, 2023). Since these datasets only have categorical metadata, we represented each categorical label by a fixed embedding. This fixed embedding is added to the output of each layer in the generator after the batch normalization layers. Similarly, we add the fixed embedding to the output of each layer in the discriminator. To learn the conditional distribution, along with predicting whether a sample is real or fake, we also predict the logit of each categorical metadata, similar to (Odena et al., 2017). In our training experiments, we noticed that predicting the metadata category for the fake sample results in poor-quality samples. Hence, during training, we only predict the category for the real samples.
@@ A.5.1 Main GAN Baselines
A.5.2 WaveGAN Implementation Details
We train the WaveGAN model for all the datasets with a learning rate of $10^{-4}$ and store the checkpoints after every 100 epochs. We sample a $48$-dimensional random vector for the Electricity, Air Quality, and Traffic datasets and a $100$-dimensional random vector for the ECG dataset. This vector serves as the noise input to the generator. We used the PyTorch implementation (Link to the repo) and the code from (Alcaraz & Strodthoff, 2023) to implement WaveGAN. We adjusted the number of parameters in the generator and discriminator to roughly match the size of our Time Weaver models.
  • For the Air Quality dataset, the total number of trainable parameters in the GAN model is 15.2 million, and the generator has 8.51 million trainable parameters.
  • For the Traffic dataset, the total number of trainable parameters in the GAN model is 13.7 million, and the generator has 7.017 million trainable parameters.
  • For the Electricity dataset, the total number of trainable parameters in the GAN model is 13.3 million, and the generator had 7.17 million parameters.
@@ A.5.2 WaveGAN Implementation Details
We chose the checkpoints that provided lower values of J-FTSD on the test data set.
A.5.3 Pulse2Pulse GAN Implementation Details
We train the Pulse2Pulse GAN model in the same manner as WaveGAN for all the datasets, with a learning rate of $10^{-4}$, and store the checkpoints after every 100 epochs. Here, the noise input to the generator has the same dimensions as the time series sample that we want to generate. We adjusted the number of parameters in the generator and discriminator to roughly match the size of our Time Weaver models.
  • For the Air Quality and Traffic datasets, the total number of trainable parameters in the GAN model is 14.1 million, and the generator has 7.45 million trainable parameters.
  • For the Electricity dataset, the total number of trainable parameters in the GAN model is 16.9 million, and the generator has 8.4 million parameters.
@@ A.5.3 Pulse2Pulse GAN Implementation Details
A.5.4 Additional GAN Baselines
In addition to WaveGAN (Donahue et al., 2019) and Pulse2Pulse GAN (Thambawita et al., 2021) models, we experimented with TTS-GAN (Li et al., 2022) and the well-established TimeGAN (Yoon et al., 2019) models. Unfortunately, we were unable to train these models to generate quality samples. These models were likely challenged by higher input time series lengths than their original implementation. For example, TimeGAN and TTS-GAN consider time steps up to 24 and 188, respectively, while we consider time steps up to 1000. A similar problem was also faced by Alcaraz & Strodthoff (2023). We include our training examples after 10000 epochs for the Air Quality and Traffic datasets in Fig. 9.
Figure 9: TimeGAN and TTS-GAN failed to generate realistic samples after 10000 epochs. This figure shows the samples generated from the test dataset after 100, 1000, and 10000 training epochs, where row 1 and row 2 correspond to TTS-GAN and TimeGAN, respectively. We can see that both models fail to generate high-quality realistic samples.
A.6 Evaluation Metrics
@@ A.6.1 J-FTSD Details
For the Electricity, Air Quality, and Traffic datasets, the horizon is 96, i.e., $L=96$. So, we consider time series and metadata patches of length $L_{\textnormal{patch}}=64$. Consequently, we obtain the time series and the metadata embeddings for these patches using $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$, respectively. We compute the J-FTSD metric from these embeddings using Eq. 6. For the ECG dataset, since the horizon is 1000, $L_{\textnormal{patch}}$ is set to 256.
One important detail is that the feature extractors, $\phi_{\textnormal{time}}$, and $\phi_{\textnormal{meta}}$, are trained on the entire dataset. This was done to ensure that the feature extractors could effectively obtain the metadata-specific features, improving the evaluation process. For example, if the feature extractors were trained only on the train and validation sets, they would not be able to detect unique trends in the test set.
A.6.2 Train on Synthetic Test on Real (TSTR) Metric Description
@@
For TSTR, we use a standard ResNet 1D (He et al., 2016) architecture. We pick the most physically relevant categorical metadata category for the classification task for each dataset. To this end, we performed the following classification tasks:
@@
  • Classification over 12 months in the Electricity dataset. We trained the classifier with cross-entropy loss for 30 epochs with a learning rate of $10^{-4}$.
@@
  • Classification over 71 heart disease statements in the ECG dataset. We note that for a given time series sample, more than one class could be active. So, we trained a classifier with binary cross-entropy loss for 200 epochs with a learning rate of $10^{-4}$.
@@
  • Classification over 11 coarse weather descriptions in the Traffic dataset. We treat the classification task here as a multi-class, multi-label classification problem. So, we trained a classifier with binary cross-entropy loss for 250 epochs with a learning rate of $10^{-4}$.
@@
  • Classification over 12 weather stations for the Air Quality dataset. We used the cross-entropy loss for 500 epochs with a learning rate of $10^{-4}$.
@@
Here, we note that with the pre-trained diffusion model, we generated the synthetic train, validation, and test datasets. The classifier is trained on the synthetic train dataset, and the checkpoints are stored with the synthetic validation dataset. We finally evaluated the model on the real test dataset. We use the Area Under the Receiver Operating Characteristic Curve (ROC AUC) as the TSTR metric.
A.7 Additional Quantitative Results
This section provides additional quantitative results and ablation studies for Time Weaver and J-FTSD.
A.7.1 Performance Comparison Between Time Weaver And GAN Models In The Presence And Absence Of Metadata
We perform an ablation study to compare the performance of Time Weaver and GAN in the presence and absence of metadata for the Air Quality and Electricity datasets. We observe the FTSD metric (Frechet Distance without the metadata embedding) for Time Weaver-CSDI and WaveGAN (Donahue et al., 2019). The FTSD metric measures the real and generated time series datasets’ distributional similarity. First, we note that in the absence of metadata, WaveGAN provides lower FTSD for the Air Quality dataset, and Time Weaver-CSDI provides lower FTSD for the Electricity dataset. However, in the presence of metadata, Time Weaver-CSDI outperforms WaveGAN in both datasets. The quantitative results are tabulated in Table 8. This result aligns with our key insight that GAN s perform poorly with continuous conditions. However, diffusion-based approaches such as Time Weaver can handle any arbitrary combination of categorical, continuous, and time-varying metadata conditions.
Approach Air Quality Electricity
Without
Metadata
With
Metadata
Without
Metadata
With
Metadata
[…]
Table 8: Time Weaver outperforms GAN s for metadata-specific time series generation for complex metadata. Here, we show a quantitative comparison between the performance of Time Weaver and GAN models in the presence and absence of metadata. We use the FTSD metric (Frechet Distance without metadata embeddings) for this comparison. We evaluate Time Weaver-CSDI and WaveGAN (Donahue et al., 2019) on the Air Quality and Electricity datasets. The key observation is that in the presence of metadata, Time Weaver-CSDI significantly outperforms WaveGAN (Donahue et al., 2019) on the FTSD metric. This indicates Time Weaver’s ability to handle any arbitrary combination of categorical, continuous, and time-varying metadata.
A.7.2 Performance Comparison Between Time Weaver And Diffusion Model With U-Net 1D Denoiser
In addition to GAN s, we compare the performance of Time Weaver against diffusion models with U-Net 1D (Ronneberger et al., 2015) as the denoiser backbone. We choose U-Net 1D as U-Net is the most commonly used denoiser in image and video synthesis. This comparison aims to highlight the requirement of time series-specific modifications to the denoiser to obtain high-quality samples. Table 9 shows the quantitative comparison between U-Net 1D and the best-performing Time Weaver model for all datasets. We choose the best-performing Time Weaver model using the TSTR metric in Table 2. Overall, we observe that Time Weaver significantly outperforms the U-Net 1D baseline by around 9% on the TSTR metric on average. Similarly, Time Weaver outperforms the U-Net 1D baseline by around 17% on the J-FTSD metric.
TSTR ($\uparrow$)
Approach Air Quality ECG Traffic Electricity
[…]
J-FTSD ($\downarrow$)
Air Quality ECG Traffic Electricity
[…]
Table 9: Time Weaver outperforms the diffusion model with a U-Net 1D denoiser backbone. Here, we show quantitative results for the comparison between the best-performing Time Weaver model and the diffusion model baseline with U-Net 1D as the denoiser backbone. Overall, we observe that Time Weaver significantly outperforms the U-Net 1D baseline by around 9% on the TSTR metric and around 17% on the J-FTSD metric on average. This superior performance highlights the necessity for time series-specific changes to the denoiser backbone in diffusion models. For example, Time Weaver-CSDI contains feature and temporal transformer layers that can learn the required correlation between the input metadata and the generated time series channels, as well as the correlation between different time stamps. Meanwhile, U-Net, a popular denoiser used for image and video generation, cannot effectively capture the temporal and channel-wise correlation between the metadata and time series.
A.7.3 Correlation Between Our Proposed J-FTSD Metric And The Dynamic Time Warping Metric
Dynamic Time Warping (DTW) (Itakura, 1975) is a similarity metric between two different time series samples. Dynamic Time Warping looks for the temporal alignment that minimizes the Euclidean distance between two aligned time series samples. Lower values of DTW indicate high similarity between the two time series samples. As a metric to evaluate metadata-specific time series generation models, DTW works effectively when there is only one time series per metadata condition. In this case, we can compare the single real time series with the generated time series for the same metadata to check if the metadata-specific features are retained. If there are multiple time series per metadata condition (which is the case in the ECG dataset), DTW might not be the most effective solution. In our experiments, we note that there is only one time series sample per metadata condition in the Air Quality, Electricity, and Traffic datasets. For these datasets, we observe that, among different approaches used in our experiments, lower values of J-FTSD correspond to lower values of DTW. The quantitative results are provided in Table 10. This further affirms the validity of our proposed J-FTSD metric.
Approach Air Quality Traffic Electricity
[…]
Table 10: Our proposed J-FTSD metric correlates with the Dynamic Time Warping metric (Itakura, 1975). This table shows the Dynamic Time Warping (DTW) metric computed between the real and generated time series samples for the Air Quality, Electricity, and Traffic datasets. Note that for the DTW metric computed between real and generated time series samples, lower values indicate higher similarity between the two samples. Comparing with the J-FTSD values in Table 2, we note that lower values of J-FTSD correspond to lower values of DTW. This further affirms the validity of our proposed J-FTSD metric. Additionally, note that our proposed Time Weaver models provide the lowest DTW for all the datasets.
A.8 Training and Inference Complexity
In this section, we provide a quantitative analysis of the training and inference complexities of our proposed Time Weaver models. For training complexity, we show how the J-FTSD metric varies over epochs during training in Fig. 10. We note that Time Weaver provides a faster rate of decrease of J-FTSD when compared to GAN s.
Figure 10: Time Weaver provides a better rate of decrease of J-FTSD during training than GAN s. Here, we show how J-FTSD changes over epochs during the training phase for the Air Quality dataset. For this experiment, we compare Time Weaver-CSDI against WaveGAN. We observe that Time Weaver-CSDI provides a much higher rate of decrease of J-FTSD when compared to WaveGAN, indicating better training.
For the inference complexity, we provide a theoretical discussion on the time complexity of Time Weaver in Sec. 6. Additionally, we provide the measured inference latency for Time Weaver variants in Table 11. From Table 11, it can be seen that the inference latency for Time Weaver-CSDI is much lower than the inference latency for Time Weaver-SSSD for generating a single sample. However, we note that Time Weaver-SSSD is more suitable for batched generation as Time Weaver-CSDI drastically limits the maximum allowable batch size due to the forward pass through its feature transformer layers.
Inference Latency (s)
Air Quality ECG Traffic Electricity
[…]
Table 11: Inference latency of Time Weaver models for generating one sample. In this table, we show the inference latency in seconds for our proposed Time Weaver models for all datasets. We compute the mean inference latency and the standard deviation for generating one sample over ten runs. The inference experiments were performed on a single NVIDIA RTX A5000 GPU.
A.9 Discussion On The Causal Relationship Between Time Series And Metadata
In this section, we present an interesting observation relevant to Time Weaver’s ability to retain the causal effects of the input metadata on the generated time series. We consider the Air Quality dataset, where physical models exist that indicate the effect of rainfall (metadata) on the particulate matter levels (one of the time series channels). We show that Time Weaver generates particulate matter levels that adhere to this effect through qualitative examples in Fig. 11. However, we note that this is not rigorous proof of causality.
Figure 11: Time Weaver retains causal relationships between the input metadata and the generated time series. In this figure, we showcase Time Weaver’s ability to retain essential causal relationships between metadata and generated time series for the Air Quality dataset. In particular, we consider the effects of rainfall (metadata) on the particulate matter levels (PM2.5, one of the channels in the generated time series). We show two instances from the dataset with a sudden spike in rainfall. In both rows, the left image shows the rainfall level metadata and the center image shows the PM2.5 time series corresponding to the rainfall metadata from the dataset. The right image shows the generated PM2.5 time series channel for the rainfall metadata input. Observe that the sudden spike in the particulate matter level is faithfully replicated in the generated samples.
A.10 Discussion On The Feature Extractors Used In The J-FTSD Metric
In this section, we analyze the embeddings from the feature extractors used to compute J-FTSD. Note that these feature extractors ($\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$) are learned using a contrastive learning-based training approach similar to CLIP (Radford et al., 2021). The feature extractors are jointly trained (check Alg. 1) to maximize the similarity between a time series embedding and its paired metadata embedding. This results in the clustering of time series embeddings for time series samples that have similar paired metadata. We test this observation on the ECG and the Air Quality datasets and show low-dimensional visualization of the time series embeddings in Fig. 12.
We use t-SNE on the time series embeddings for dimensionality reduction. For the ECG dataset, we observe that time series samples corresponding to the most common heart disease statement get clustered together. This is shown in Fig. 12. Similarly, for the Air Quality dataset, we observe that the time series samples get clustered into 4 clusters. However, the dataset has a combination of categorical metadata conditions such as stations, years, months, etc, and continuous metadata conditions such as temperature, pressure, rain levels, etc. This complexity in metadata conditions inhibits us from interpreting the clusters, as in the case of the ECG dataset. Therefore, our critical insight is that the contrastive learning approach results in the clustering of the time series embeddings corresponding to similar paired metadata. However, the interpretability of the clusters is non-trivial for arbitrary combinations of categorical, continuous, and time-varying metadata conditions.
Figure 12: Interpretability of the time series embeddings obtained from J-FTSD’s time series encoder. In this figure, we show the clustering of time series embeddings obtained from the time series feature extractor $\phi_{\textnormal{time}}$. We use t-SNE on the time series embeddings for dimensionality reduction. Note that the feature extractors ($\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$) are trained jointly using a contrastive learning-based approach (check Alg. 1). We show results for clustering on two datasets - ECG and Air Quality. For the ECG dataset, we showcase that the samples corresponding to the most common disease get clustered together. For the Air Quality dataset, we show that the time series embeddings get clustered into 4 clusters. However, due to the complexity of the metadata conditions (a combination of categorical and time-varying continuous conditions), we cannot interpret the clusters as in the case of the ECG dataset.
A.11 Additional Qualitative Results
@@
Figure 13: Generated time series samples from the Electricity dataset
Figure 14: Qualitative Results from the Time Weaver-CSDI model for the Traffic dataset
Figure 15: Qualitative results from the Time Weaver-CSDI model for the Air Quality dataset